### v2

# Direct determination of atomic number density in MEMS vapor cells via single-pass absorption spectroscopy (SPAS)

Abstract
Micro-electro-mechanical systems (MEMS)-based chip-scale alkali vapor cells are the essential components in emerging quantum technologies, including compact atomic clocks, chip-scale magnetometers, and miniature quantum opto-electronic systems. The sensitivity of MEMS-based atomic quantum technology devices depends on the atomic number density. Thus, it is important to have an accurate estimate of the atomic number density in chip-scale alkali vapor cells to optimize light-matter interactions and design efficient quantum sensing systems. Here, we present a direct and quantitatively validated method for determining the rubidium (Rb) number density in warm alkali vapor using single-pass absorption spectroscopy (SPAS). The absolute transmission spectra are measured and modeled using the 780.24 nm as well as the 420.29 nm transition in a Rb-filled MEMS vapor cell. The atomic number density measurements and the model were also validated using a commercial vapor cell of length 100 mm. The theoretical model employs a density-matrix formalism within the Lindblad framework and incorporates directly measurable experimental parameters, such as laser beam power, diameter, and cell temperature. The model explicitly accounts for optical pumping, Doppler broadening, and transit-time broadening effects and exhibits quantitative agreement ($99\%$) with experimental spectra over a broad range of temperatures (293–343 K), laser intensities ($\sim 0.2\,I_{\mathrm{sat}}$ to $\sim 2\,I_{\mathrm{sat}}$), and cell lengths (2–100 mm). The extracted densities from the MEMS cell closely follow the empirical vapor-pressure model by Alcock et al. The demonstrated methodology provides a practical, well-controlled method for determining the atomic number density in alkali vapor cells relevant to the characterization and development of compact alkali-vapor-based devices for quantum sensing and metrology.
Single-Pass Absorption Spectroscopy, number density, Density matrix, MEMS
preprint: APS/123-QED
I Introduction
With the growing demand for portable quantum devices for use in precision timing, sensitive magnetometry [1], electric-field sensing [2, 3, 4] , and entangled photon-pair generation [5, 6], etc., micro-fabricated chip-scale vapor cells based on micro-electro-mechanical systems (MEMS) technology [7] have become an indispensable tool in the domain of field-deployable atomic quantum technology devices for use in quantum sensing and quantum communication applications. These miniaturized vapor cells enable wide applications in quantum technologies, including compact atomic clocks [8], high-precision magnetometers [9], frequency-metrology platforms [10], and integrated photonic–atomic sensors [11]. Their compact form factor, low power consumption, greater reliability, and low cost for scalable fabrication make them particularly attractive for next-generation quantum applications. Most quantum devices rely on the well-defined optical and spectroscopic properties of alkali vapors, enabling applications ranging from inertial navigation [12] and timing synchronization [13] to biomedical applications through ultra-low magnetic field detection [14].
In chip-scale quantum devices, particularly MEMS-based vapor cells, the intrinsically short optical path length makes the optical depth extremely sensitive to the atomic number density $N$. Since the stability and sensitivity of many atomic and quantum technologies scale as $1/\sqrt{\mathcal{N}}$, where $\mathcal{N}$ is the effective number of atoms/ions participating in the light–matter interaction [15, 16], precise knowledge of the ensemble density is essential for accurate device design for optimized performance and quantitative modeling. Reliable determination of number density is therefore critical for optimizing the performance of miniature atomic clocks, optically pumped magnetometers, photon-pair sources, nonlinear optical interfaces, and quantum memory systems, while avoiding density-dependent systematic shifts and unwanted broadening effects.
In alkali atoms, rubidium (Rb) is one of the most widely used platforms for MEMS-based vapor-cell applications due to its well-defined spectroscopic structure, strong optical transitions, and accessible vapor pressures near room temperature [17]. Naturally occurring Rb has two stable isotopes, 85Rb (72.2%) [18] and 87Rb (27.8%) [19], each exhibiting distinct multi-level hyperfine structures that must be accurately modeled to extract quantitative information of the atomic number density. Estimating the number density of warm alkali vapors typically relies on empirical vapor-pressure relations [20]. The standard expression for the temperature-dependent vapor pressure $P_{\mathrm{vap}}(T)$ was first proposed by Killian [21] and later refined by Alcock, Itkin, and Horrigan [20]. For rubidium, the vapor pressure $P_{\mathrm{vap}}(T)$ (in torr) as a function of temperature $T$ (in Kelvin) is given by
$\log_{10}P_{\mathrm{vap}}(T)=2.881+a-\frac{b}{T},$ (1)
where the constants $a$ and $b$ correspond to different physical phases: $a=4.857,b=4215~\text{(solid phase)},a=4.312,b=4040~\text{(liquid phase)}.$ Here, $P_{\mathrm{vap}}$ is expressed in torr; when expressed in atmospheres, the additive constant 2.881 is omitted. The corresponding atomic number density $N(T)$ is determined from the ideal-gas relation,
$N(T)=\frac{P_{\mathrm{vap}}(T)}{k_{\mathrm{B}}T},$ (2)
where $k_{\mathrm{B}}$ is the Boltzmann constant and $P_{\mathrm{vap}}$ is converted to Pascals using $1~\text{torr}=133.322~\text{Pa}$. This empirical formulation remains the standard benchmark for validating experimentally inferred alkali-vapor densities.
Various experimental techniques have been employed to determine the atomic number density in warm alkali vapor cells, including fluorescence spectroscopy [22], Faraday rotation [23], and spin-exchange measurements [24]. In fluorescence spectroscopy, the population of the excited state can be inferred from the intensity of the emitted light. In contrast, Faraday rotation correlates the polarization rotation of a probe beam with resonant atom density. Measurement of the number density based on the Faraday rotation technique relies on the spin polarization of the atomic sample. It requires sophisticated magnetic-field isolation using mu-metal shields, as well as precision electronics for detecting spin polarization with balanced photodetectors. Thus, implementing this method of number density measurement in field environments outside the laboratory is challenging. The spin-exchange method infers the number density from collisional relaxation rates. Still, it is usually limited to specific regimes, e.g., high optical depth at substantially elevated temperatures, where direct absorption is impractical. Despite the utility of these approaches, single-pass absorption spectroscopy (SPAS) [25] remains one of the most attractive, simple, and widely accessible methods for characterizing warm vapor systems, especially near room temperature [26]. However, while SPAS has been extensively used to investigate spectral lineshapes [27] and transition strengths [26], it has rarely been used for the explicit, quantitative determination of absolute number density in a MEMS-based vapor cell. This limitation arises from the need for accurate baseline calibration and the complexities introduced by optical pumping, multi-level dynamics, and Doppler broadening in chip-scale atomic vapor cells.
In this work, we present a systematic, quantitatively validated framework for determining the atomic number density of a dilute atomic vapor using a multi-level master equation model. Our theoretical approach efficiently incorporates the intrinsic differences between the two stable isotopes, 85Rb and 87Rb, enabling accurate and computationally tractable modeling of their spectroscopic signatures. By treating both isotopes within a unified formalism, the model not only facilitates direct comparison with experimental data but also offers insights into the role of hyperfine structure, optical pumping, and relaxation mechanisms, such as transit-time broadening and finite laser linewidth. Importantly, our approach does not rely on conventional curve fitting of arbitrary parameters; instead, we incorporate all experimentally measured quantities, such as cell length, laser power, beam diameter, and temperature, into a theoretical model based on Lindblad formalism and extract the atomic number density as the only free parameter by matching the calculated absorption spectrum to the measured data. To the best of our knowledge, the work reported here represents the first measurement of the total number density of a dilute alkali vapor using single-pass absorption spectroscopy across a wide range of temperatures, laser beam powers, and different absorption lengths.
We demonstrate this method across markedly different optical path lengths, using both a MEMS vapor cell (2 mm, designed and fabricated in the LEOS (Laboratory for Electro-Optics Systems), ISRO (Indian Space Research Organization)[28]) and a conventional 100 mm (from Triad Technology Inc.) vapor cell, and at two widely separated wavelengths corresponding to the $5S_{1/2}\rightarrow 5P_{3/2}$ (780.24 nm) and $5S_{1/2}\rightarrow 6P_{3/2}$ (420.29 nm) transitions. This broad validation highlights the generality of the technique and its relevance for precision quantum-optics experiments, spectroscopy in weak-absorption regimes, and the characterization of chip-scale alkali-vapor devices. Although our study focuses on Rb, the approach is readily extendable to other atomic and molecular species, making it a versatile tool for laboratory and field-deployable sensing applications.
The manuscript is organized as follows. In section II, we outline the theoretical framework for modeling the absorption spectra, detailing the development of a multi-level master equation approach and integrating Doppler broadening into absorption coefficient calculations. In section III we describe the experimental methodology, including the layout of the optical setup and the techniques used for signal acquisition. In the following section section IV, we present a comparative analysis of the simulated and experimentally observed spectra, emphasizing the influence of critical parameters, such as temperature and laser power, on the number density estimation. Finally, in section V, we summarize key findings and discuss potential avenues to extend this approach to other atomic and molecular systems for sensing and metrology applications.
II Theoretical background
In the presence of light–matter interaction and dissipative processes such as dephasing and spontaneous emission, the state vector alone is no longer sufficient to capture the complete dynamics of the system [29]. Instead, the density matrix denoted by $\rho$ is used because it can describe both pure and mixed states, including decoherence, spontaneous emission, and other dissipative effects [30, 31, 32]. For an $N$-level quantum system, the density matrix is an $N\times N$ Hermitian matrix. In this matrix, the diagonal elements $\rho_{ii}$ indicate the probability (or population) of finding the system in the $i$-th state, and the off-diagonal elements $\rho_{ij}$ represent the quantum coherence between the states $|i\rangle$ and $|j\rangle$. This matrix framework is essential in scenarios where a quantum system interacts with external fields or environments, leading to coherence loss.
Figure 1: Energy-level diagram of rubidium (87Rb) showing the 420 nm excitation scheme, relevant intermediate states, and decay pathways, highlighting the 780 nm and 420 nm transitions with their corresponding branching ratios [33] used to determine the atomic number density of the Rb atoms in the MEMS vapor cell.
In this work, we use the density-matrix formalism, in which the atom is treated quantum-mechanically, and the electric field is treated classically. We construct the Hamiltonian of a four-level system for the transition of Rb atom from ground state $5S_{1/2}$ to the excited state $5P_{3/2},6P_{3/2}$ by including the relevant atomic energy levels (as shown in Figure 1) and their interactions with the light field. We solve the steady state of the system using the master equation. The results accurately depict the coherence terms, including both absorption and dispersion. To simulate the absorption spectroscopy of these Rb transitions ($5S_{1/2}\to 5P_{3/2},6P_{3/2}$), we construct a theoretical model using two separate four-level atomic systems, for each isotope: $^{85}\mathrm{Rb}$ and $^{87}\mathrm{Rb}$. Each system includes a ground state and three excited states. For a given isotope, the model distinguishes between transitions (dipole allowed) originating from the lower ($|0_{a}\rangle$) and upper ($|0_{b}\rangle$) hyperfine ground states, which couple to the common excited states $|2\rangle$ and $|3\rangle$, and to the exclusive excited states $|1\rangle$ and $|4\rangle$, respectively, as shown in Figure 2(b). This framework enables accurate modeling of optical pumping effects and decay pathways, incorporating the branching ratios between states.
Figure 2: (a) [Left] Generalized four-level atomic system comprising a ground state $|0\rangle$ and three excited states $|1\rangle$, $|2\rangle$, and $|3\rangle$. Transitions are driven by classical fields with Rabi frequencies $\Omega_{01}$ , $\Omega_{02}$ , and $\Omega_{03}$ , and corresponding detunings $\Delta_{1}$ , $\Delta_{2}$ , and $\Delta_{3}$ . Spontaneous emission processes occur with decay rates $\Gamma_{10}$ , $\Gamma_{20}$, and $\Gamma_{30}$ . (b) [Right] Schematic energy level diagram representing the hyperfine structure of rubidium atoms. Transitions originate from two ground hyperfine states, $|0_{a}\rangle$ and $|0_{b}\rangle$, coupling to excited states $|1\rangle$ to $|4\rangle$ as allowed by dipole selection rules. States $|2\rangle$ and $|3\rangle$ are common to both pathways. Decay channels follow branching ratios defined by $\Gamma_{b}$ and $\Gamma_{d}$ (from $|3\rangle$) and $\Gamma_{c}$ and $\Gamma_{e}$ (from $|2\rangle$).
[…]
To derive such a Hamiltonian [34, 35]. The interaction Hamiltonian has the general form $H_{I}(t)=-\textbf{{d}}\cdot\textbf{{E}(t)}$, where d is the electric dipole operator and E(t) is the electric field. Expand the dipole operator in the atomic basis and introduce the Rabi frequency $\Omega_{ij}=\frac{E_{0}}{\hbar}D_{ij}$ [36], which quantifies the strength of the coupling between the ground $|i\rangle$ and excited states $|j\rangle$ (by definition $\Omega_{ij}=\Omega_{ji}^{*}$ and $\Omega_{ii}=0$), where $E_{0}$ is the field amplitude and $D_{ij}$ is the dipole matrix element [37]. By applying the rotating-wave approximation (RWA) to keep only the absorption $|i\rangle\to|j\rangle$ and emission $|j\rangle\to|i\rangle$ terms, and then transforming into the rotating frame to get the time-independent Hamiltonian of the four-level system,
$H=\hbar\begin{bmatrix}0&\frac{{\Omega}_{01}}{2}&\frac{\Omega_{02}}{2}&\frac{\Omega_{03}}{2}\\ \frac{\Omega_{10}^{*}}{2}&-\Delta_{1}&0&0\\ \frac{\Omega_{20}^{*}}{2}&0&-\Delta_{2}&0\\ \frac{\Omega_{30}^{*}}{2}&0&0&-\Delta_{3}\end{bmatrix}$, (3)
where $\Omega_{01}$, $\Omega_{02}$ and $\Omega_{03}$ are the Rabi frequencies associated with the three transitions, and $\Delta_{1}$, $\Delta_{2}$ and $\Delta_{3}$ are the corresponding detunings Figure 2(a).
II.1 Master equation for density matrix evolution
To realistically describe the dynamics of an atomic system interacting with external laser fields and subject to environmental decoherence, one must employ an open quantum system formalism. In this context, the evolution of the system’s density matrix $\rho$} is governed by the Lindblad master equation [38, 30]:
$\frac{\partial{\rho}}{\partial t}=-\frac{i}{\hbar}[{H},{\rho}]+{\mathcal{L}}({\rho}),$ (4)
where H{H} is the total Hamiltonian of the system, and the term ${\mathcal{L}}({\rho})$ represents the non-unitary contribution due to dissipative processes, including both decay and dephasing terms. This formulation enables the inclusion of essential effects such as spontaneous emission, dephasing, and population relaxation.
The general form of the Lindblad superoperator acting on the density matrix is given by [30, 39]:
${\mathcal{L}}({\rho})=\sum_{k}\left({C}_{k}{\rho}{C}_{k}^{\dagger}-\frac{1}{2}\{{C}_{k}^{\dagger}{C}_{k},{\rho}\}\right),$ (5)
where ${C}_{k}$ denotes the collapse (or jump) operators that model the interaction between the system and its environment. For pure dephasing of an excited state $|e\rangle$, the collapse operator is defined as ${C}_{e}=\sqrt{\gamma_{e}}|e\rangle\langle e|$, with $\gamma_{e}$ representing the pure dephasing rate associated with state $|e\rangle$. In contrast, spontaneous emission from an excited state $|e_{j}\rangle$ to a ground state $|g_{i}\rangle$ is described by the collapse operator as, ${C}_{ij}=\sqrt{\Gamma_{ij}}|g_{i}\rangle\langle e_{j}|$, where $\Gamma_{ij}$ denotes the decay rate from $|e_{j}\rangle$ to $|g_{i}\rangle$.
II.2 Electrical susceptibility and absorption coefficient
In the theoretical model, the interaction between the atomic ensemble and the optical fields are described by the density matrix $\rho$, and the evolution is governed by the master equation given in Equation 4. For continuous-wave excitation, the atomic system is driven into a time-independent equilibrium due to relaxation and decoherence processes such as spontaneous emission, transit-time decay, and dephasing. Consequently, the steady-state condition $\partial\rho/\partial t=0$ [34] is sufficient to describe the dynamics of the system. The steady-state solution of the density matrix gives the coherence terms $\rho_{ij}$ corresponding to the optical transitions, from which the complex electric susceptibility of the medium can be obtained, which characterizes its polarization response to an external electromagnetic field. For a dipole-allowed transition between states $|i\rangle\to|j\rangle$, the complex susceptibility as a function of detuning ($\Delta,T$) can be expressed as [34],
$\chi_{ij}(\Delta,T)=-\frac{2N(T)D_{ji}^{2}}{\hbar\epsilon_{0}\Omega_{ij}}\rho_{ij}(\Delta),$ (6)
where $N$ is the atomic number density [20], $\Omega_{ij}$ is the Rabi frequency of the probe field describing the light–matter coupling strength, and $\rho_{ij}(\Delta)$ denotes the optical coherence obtained from the steady-state solution of the density-matrix equations (see Equation 4). The dipole matrix element is given by,
$D_{ji}=\langle j|D|i\rangle=\sqrt{3}\sqrt{\frac{3\varepsilon_{0}h\lambda^{3}\Gamma_{ji}}{8\pi^{2}}},$ (7)
with $\lambda$ the transition wavelength and $\Gamma_{ji}$ the effective decay rate. The latter is expressed as $\Gamma_{ji}=C_{f}^{2}\times\mathrm{\textit{BR}_{ji}}\times\Gamma$, where $\Gamma$ is the natural decay rate, $\mathrm{\textit{BR}_{ji}}$ the branching ratio of the transition $|j\rangle\to|i\rangle$, and $C_{f}$ the Clebsch–Gordan coefficient associated with the dipole coupling [26, 40].
In our theoretical model, we explicitly consider the hyperfine structure of the rubidium atom at 780 nm and 420 nm ($5^{2}S_{1/2}\leftrightarrow 5^{2}P_{3/2},6^{2}P_{3/2}$ ) transitions, constructing two distinct four-level systems for the two naturally occurring isotopes, 87Rb and 85Rb. As shown in Figure 2(b) for 87Rb, the hyperfine levels of the ground state are $|0_{a}\rangle=|F=1\rangle$ and $|0_{b}\rangle=|F=2\rangle$, while the hyperfine excited state ($5^{2}P_{3/2},6^{2}P_{3/2}$) consists of $\lvert 1\rangle$ to $\lvert 4\rangle$, corresponding to $\lvert F^{\prime}=0\rangle$ to $\lvert F^{\prime}=3\rangle$ respectively. For 85Rb, the ground state levels are $|0_{a}\rangle=\lvert F=2\rangle$ and $\lvert 0_{b}\rangle=\lvert F=3\rangle$ , with excited state levels $\lvert 1\rangle$ to $\lvert 4\ranglecorresponding$ to $\lvert F^{\prime}=1\rangle$ to $\lvert F^{\prime}=4\rangle$ respectively.
[…]
$\text{Branching Ratio}=BR_{if}=\frac{C_{f}{i\rightarrow f}}{\sum_{i^{\prime}}C_{f}{i^{\prime}\rightarrow f}},$ (8)
where $C_{f}{i\rightarrow f}$ is the transition strength from the initial state $\lvert i\rangle$ to a particular final state $\lvert f\rangle$. Since the total electronic angular momentum $J$, nuclear spin $I$, and orbital angular momentum $L$ are identical for the $5P_{3/2}$ and $6P_{3/2}$ states, the hyperfine structure, Clebsch–Gordan coefficients, and branching ratios for individual hyperfine transitions remain unchanged relative to the $D_{2}$ line. Despite these similarities, the overall absorption strength of the $5S_{1/2}\rightarrow 6P_{3/2}$ transition is significantly weaker than that of the $D_{2}$ transition. Unlike the $D_{2}$ line, where spontaneous decay from the excited state occurs exclusively to the ground state, the $6P_{3/2}$ state exhibits multiple radiative decay channels, as shown in Figure 1. Approximately $23\%$ of the total spontaneous decay from $6P_{3/2}$ returns directly to the ground state $5S_{1/2}$ [33], with the remaining population decaying via intermediate states. This reduced branching ratio to the ground state and a smaller dipole matrix element led to substantially weaker absorption on the $420~\mathrm{nm}$ transition. The branching ratios corresponding to all allowed transitions of both 85Rb and 87Rb are presented in Table 1.
$F_{g}$ $F_{e}$
1 2 3 4
2 1 79\ $\frac${7}{9}\ $\frac{4}{9}$ NA
3 NA $\frac{2}{9}$ $\frac{5}{9}$ 1
$F_{g}$ $F_{e}$
0 1 2 3
1 1 $\frac{1}{2}$ $\frac{1}{6}$ NA
2 NA $\frac{1}{2}$ $\frac{5}{6}$ 1
Table 1: Branching ratio for the $D_{2}$ line of (a) 85Rb and (b) 87Rb, where ’NA’ is not allowed as per the selection rule.
The absorption coefficient follows from the imaginary part of the susceptibility,
$\alpha_{ij}(\Delta,T)=k*\mathrm{Im}{\chi_{ij}(\Delta,T)},$ (9)
where $k$ is the probe wavevector. In the four-level model considered here (see Figure 2(a)), the ground state couples to three excited states, giving rise to three distinct absorption channels $\alpha_{0j}$ ($j=1,2,3$).
II.3 Transit time and Doppler broadening
The linewidth of an absorption spectrum in an atomic vapor cell is influenced by several homogeneous and inhomogeneous broadening mechanisms, among which transit-time relaxation and Doppler broadening play significant roles. The transit-time relaxation arises from the interaction time between moving atoms and the incident laser beam. As atoms with velocities from the Maxwell-Boltzmann distribution traverse the probe region, they are effectively reset upon entering and exiting. When an atom leaves the interaction volume, it is statistically replaced by a new, unpumped atom from the surrounding ensemble. From the system’s perspective, this process acts as an additional relaxation pathway that contributes to linewidth broadening. Unlike spontaneous emission, which occurs only from excited states to the ground state, transit-time relaxation introduces a dephasing mechanism that affects both ground and excited states uniformly. It destroys the phase coherence of the atomic ensemble, thereby broadening the spectral line. The transit relaxation rate, denoted as $\gamma_{t}$ expressed as,
$\gamma_{t}=\frac{\langle v\rangle}{D},$ (10)
where $\langle v\rangle=\sqrt{\frac{8k_{B}T}{\pi m}}$ is the mean thermal velocity of the atoms, $k_{B}$ is the Boltzmann constant, $T$ is the absolute temperature of the vapor, $m$ is the atomic mass, and $D$ is the FWHM of the probe beam [43]. Thus, a smaller beam diameter or a higher atomic temperature leads to a larger $\gamma_{t}$, corresponding to a stronger transit-induced broadening. The transit time broadening contributes an additional dephasing rate [44] and is added to the dephasing part of the Lindblad superoperator Equation 5.
In a warm vapor, the atomic motion gives rise to Doppler broadening, an inhomogeneous effect arising from the Maxwell–Boltzmann distribution of atomic velocities. At finite temperatures, atoms move with varying velocities according to the Maxwell–Boltzmann distribution, causing each atom to experience a different effective laser frequency due to the Doppler shift. As the temperature increases, the distribution widens, resulting in a broader spectral line. The absorption coefficient derived in Equation 9 describes the natural line shape of an atomic transition, which follows a Lorentzian profile characterized by a homogeneous linewidth that includes contributions from spontaneous emission and transit-time relaxation (see Equation 10). Considering motion only along the laser propagation direction (z-axis), the Doppler-shifted detuning between the laser frequency and the atomic resonance is $\Delta_{\text{modified}}=\Delta\mp kv_{z}$, where the sign depends on the direction of motion of the atoms, $k$ is the wavevector, and $v_{z}$ is the velocity component along the $z$-axis. The overall observed line shape is obtained by convolving the Lorentzian profile with the Maxwell–Boltzmann velocity distribution $M(v)$, which accounts for the spread in velocity along the propagation direction. This convolution yields the Voigt profile, a line shape that accurately represents the combined effects of homogeneous (natural and transit-time) and inhomogeneous (Doppler) broadening mechanisms in thermal atomic vapors.
To obtain the Doppler-broadened spectra, the absorption coefficient from Equation 9 is integrated over all atomic velocities,
$\alpha^{\prime}_{ij}(\Delta,T)=\int_{-\infty}^{+\infty}\alpha_{ij}(\Delta-kv,T)M(v)dv.$ (11)
Ideally, the simulation of the Doppler-broadened absorption coefficient (in Equation 11) should span a velocity range that captures nearly all atoms in the vapor. However, to balance efficiency and computation time, we consider the velocity range from $4\sigma_{v}$ to $4\sigma_{v}$, which includes approximately $99.99\%$ of the atomic population. Here, $\sigma_{v}=\sqrt{\frac{k_{B}T}{m}}$ denotes the root-mean-square (rms) velocity of the atoms along the laser propagation direction, where $k_{B}$ is the Boltzmann constant, $T$ is the temperature of the atomic vapor, and mm is the atomic mass.
II.4 Light propagation through the vapor cell
The propagation of resonant light through an atomic vapor is governed by absorption and dispersion processes that modify the transmitted intensity along the direction of propagation. For a monochromatic beam incident on a homogeneous vapor of length $z$, the transmitted intensity follows the Beer-Lambert law,
$I(z)=I_{0}e^{-\alpha^{\prime}_{ij}(\Delta,T)*z},$ (12)
where $I_{0}$ is the incident intensity, and $\alpha^{\prime}_{ij}(\Delta,T)$ is the Doppler-broadened absorption coefficient (from Equation 11), which depends on the detuning $\Delta$ of laser and the vapor temperature $T$. The absorption coefficient includes the combined effects of the atomic line profile, natural, transit time, and Doppler broadening.
Spectroscopic measurements are performed by monitoring the fraction of light transmitted through the vapor cell, and transmission provides a dimensionless and experimentally observable quantity that reflects the frequency-dependent absorption of the medium. It is therefore convenient to express the light–matter interaction in terms of the normalized transmission. The normalized transmission through the vapor cell is expressed as,
$\mathcal{T}(z)=\frac{I(z)}{I_{0}}=e^{-\alpha^{\prime}_{ij}(\Delta,T)*z}.$ (13)
III Experimental setup
We performed single-pass absorption spectroscopy [25, 26] on naturally abundant rubidium vapor contained in both a conventional vapor cell of length 100 mm and a microfabricated MEMS-based vapor cell of length 2 mm. The measurements were carried out primarily on the $D_{2}$ transition of rubidium at 780.24 nm, as shown in Figure 2. A schematic of the experimental setup is presented in Figure 3.
Figure 3: Schematic of the experimental setup for absolute absorption spectroscopy and frequency calibration. The absorption signal transmitted through the rubidium vapor cell is detected using photodetector (PD-1), while the reference transmission signal from the Fabry–Pérot interferometer (FPI) is monitored by photodetector (PD-2). Both signals are simultaneously recorded with a digital storage oscilloscope (DSO). The piezoelectric transducer (PZT) scan signal from the external cavity diode laser (ECDL) controller is used as the trigger input to synchronize the oscilloscope trace with the frequency scan. Optical components include mirrors (M1–M5), neutral density filters (NDF), lenses (L1–L3), polarizing beam splitters (PBS), an electro-optic modulator (EOM), and a beam dump (BD). The vapor cell temperature is stabilized using a temperature controller and resistive foil heaters, which are monitored by four NTC (Negative Temperature Coefficient) temperature sensors.
A tunable external-cavity diode laser (ECDL, Toptica DL PRO 780) operating at $780.24$ nm, equipped with an inbuilt $35$ dB optical isolator, was used as the light source. The output beam was spatially aligned using three-axis mirrors (M1 and M2) and subsequently divided into a spectroscopy arm and a frequency-reference arm using a half-wave plate (HWP) and a polarizing beam splitter (PBS). The optical power in the spectroscopy arm was adjusted using an additional HWP-PBS combination and a neutral density filter (NDF).
[…]
The vapor cell was enclosed in a custom-fabricated heating assembly integrated with a temperature controller (Thorlabs TC300B) to facilitate spectroscopic measurements over a range of temperatures. However, the temperature displayed by the controller does not accurately reflect the actual temperature of the vapor cell due to thermal gradients in the heating assembly. Four precalibrated NTC temperature sensors were attached at different locations in the cell to accurately determine the cell temperature (as illustrated in Figure 3). The temperature readings from these sensors were recorded using an Arduino Uno, and the average value was considered as the effective vapor cell temperature for subsequent simulations. The absorption signal was detected with an unbiased photodetector (PD-1, Thorlabs PDA36A2) and recorded on one channel of a 4-channel digital storage oscilloscope (Tektronix MSO44B).
On the other hand, for the frequency-reference measurement, the second beam was passed through an electro-optic modulator (Qubig PM7_\_NIR_\_25) driven by a $25$ MHz radio-frequency (RF) signal, which modulated the frequency of the laser beam. The modulated beam was then collimated using a lens system (L1 and L2) and directed through a mode-matching lens (L3) before entering a Fabry-Pérot interferometer (Thorlabs SA30-73 with FSR= $1500$ MHz). The transmitted signal from the interferometer, including sidebands from the 25 MHz modulation, was captured by another photodetector (PD-2) and recorded on a separate channel of the oscilloscope. This signal provides a precise reference for frequency calibration. The rubidium absorption signal and the interferometer transmission peaks were acquired simultaneously within the same oscilloscope time window, ensuring synchronized data acquisition for accurate analysis.
For quantitative analysis of absolute absorption, it is essential to calibrate the baseline of the detected transmission signal. The baseline offset arises from the intrinsic dark current of the photodetector. This offset voltage was measured independently by blocking the probe beam and optically isolating the detector from ambient light. The measured offset was subsequently subtracted from all recorded transmission signals and used to define the zero-transmission reference level. Another way to measure the zero level is to make the atomic vapor optically opaque by significantly increasing its temperature, so that probe light is not transmitted through the medium [45]. Although this technique offers a physical reference point for complete absorption, it is not universally applicable. In particular, for miniaturized vapor cells (MEMS cells) with inherently low atomic density or for investigating any weak transitions that require extreme heating to achieve complete opacity, this technique becomes difficult to implement. In such scenarios, the dark-current-based method offers a more flexible and non-invasive strategy for baseline normalization.
Non-linearities in both the laser frequency scan and the output power, introduced by the piezoelectric actuator used for frequency tuning, were corrected by the methodology prescribed in [45]. The frequency non-linearity was removed by mapping the Fabry-Pérot interferometer (FPI) transmission peaks to their known frequency separations, and the power non-linearity was removed using the off-resonant region of the measured transmitted spectrum. The off-resonant regions were fit with a higher-order polynomial to characterize the power variation across the scan, which was subsequently used to normalize the measured absorption signal.
In addition to the absolute absorption measurement at 780.24 nm, absorption spectroscopy was also performed on the $5S_{1/2}\to 6P_{3/2}$ transition at 420.29 nm using the same experimental methodology as shown in Figure 3. The optical layout and measurement protocol remained unchanged, with wavelength-appropriate substitutions for the laser source, photodetectors, and coated optical components. All data acquisition, baseline correction, and frequency calibration procedures were identical to those employed at 780.24 nm. It enables a direct comparison of number-density estimates obtained from various transitions with different strengths and saturation intensities.
IV Results and discussion
We systematically measured the single-pass absorption spectrum of rubidium vapor under various experimental conditions to determine the effects of probe power, cell temperature, and cell length. All of these parameters are precisely measured and incorporated into the theoretical model, and the model spectra were then fitted to the experimental data by treating the number density as a free parameter to achieve quantitative agreement.
The atomic number density of the rubidium vapor can be estimated theoretically from its equilibrium vapor pressure. The temperature dependence of the vapor pressure of alkali metals has been comprehensively studied, with reliable empirical relations reported by Alcock et al. [20] over a wide temperature range (298–550 K). From the experimental absorption spectra, we extracted the atomic number density by fitting the data with our theoretical model. To verify the accuracy of the measured number density, the obtained values were compared with those estimated from the empirical relation Equation 1 provided by Alcock et al. for rubidium vapor pressure and number density as a function of temperature (see Equation 2). This comparison provides a stringent consistency check of both the experimental procedure and the theoretical modeling.
Figure 4: Comparison of simulated and experimental transmission spectra of the rubidium MEMS vapor cell at $307.35$ K, $324.25$ K, and $332.25$ K. Lower panels show the residuals (experiment-simulation).
For the MEMS vapor cell (length 22~$m$) containing naturally abundant rubidium (72.2% 85Rb and 27.8% 87Rb), single-pass absorption spectra were recorded at various temperatures: ranging from 293363293-363K using a probe intensity of approximately $\sim 0.2~I_{\text{sat}}$. The corresponding experimental and theoretically fitted spectra, along with their residuals, are shown in Figure 6. The fitted spectra yield coefficients of determination $R^{2}>0.99$ for all temperatures, indicating excellent quantitative agreement. The cell temperature was monitored using precalibrated NTC temperature sensors with an intrinsic accuracy of $\pm~0.3$ K. The cell was mounted in a machined aluminum block equipped with four symmetrically placed sensors to ensure uniform heating. At elevated temperatures, a residual temperature gradient of approximately 1–2 K was observed, which accounts for the slightly larger error bars in the extracted number density at higher temperatures. As shown in Figure 6, the error bars in the MEMS cell measurements are substantially smaller, indicating the improved temperature uniformity and more precise determination of the number density. These observations emphasize that precise temperature monitoring and control are crucial for reducing uncertainty in vapor density measurements, particularly at higher operating temperatures.
We performed the same set of measurements with a commercial 100 mm Rb vapor cell with both wavelengths, 780 nm and 420 nm. Since the atomic number density is a thermodynamic property of the vapor, it is independent of the optical interaction length. The results presented in Appendix Appendix A, confirm that the same density values are obtained for both the MEMS and the standard-length cells.
Furthermore, to verify the validation of our model beyond the weak-probe regime [46], we perform the absorption spectroscopy measurement at a higher probe intensity ($\sim 2~I_{\text{sat}}$). As shown in Appendix Appendix B (see Figure 11 for more details), the model continues to reproduce the experimental spectra with $R^{2}>0.99$, confirming its robustness over a wider range of excitation conditions.
In addition to spectroscopy on the $D_{2}$ transition ($5S_{1/2}\rightarrow 5P_{3/2}$) at 780.24 nm, we also performed measurements using a $420~\mathrm{nm}$ laser addressing the $5S_{1/2}\rightarrow 6P_{3/2}$ transition. The number density is an intrinsic thermodynamic property of the rubidium vapor, determined solely by the vapor pressure at a given temperature, and is independent of the specific excited state used for optical interrogation. The agreement in the extracted number density values from both transitions thus provides a strong internal consistency check. For a given vapor cell temperature, the extracted atomic number density remains identical irrespective of whether the atoms are excited via the $5S_{1/2}\rightarrow 5P_{3/2}$ transition or the $5S_{1/2}\rightarrow 6P_{3/2}$ transition.
Figure 5: Experimental and simulated transmission spectra comparing the 780 nm and 420 nm rubidium transitions at the same optical power of $0.2\,~I_{\mathrm{sat}}$ and nearly identical temperatures (333.25 K and 332.25 K). The upper panel shows simulated spectra (solid red: 780 nm, dashed black: 420 nm) with corresponding experimental data (markers). The lower panels present the residuals.
Figure 5 compares representative transmission spectra recorded at 780 nm and 420 nm under nearly identical temperature conditions and at the same normalized probe intensity ($0.2,I_{\mathrm{sat}}$). As expected, the absorption strength of the 420 nm transition is significantly weaker than that of the $D_{2}$ transitions. The successful application of the model to MEMS cells has important implications for the development of chip-scale atomic devices, where precise knowledge of the atomic number density is crucial for optimizing sensor performance. The ability to accurately determine the vapor density in these miniaturized systems enables more accurate calibration and characterization of MEMS-based atomic, quantum, and precision timing sensors.
The extracted number density of rubidium in the MEMS cell as a function of temperature is shown in Figure 6, where results obtained from 780 nm and 420 nm absorption measurements over a range of temperatures 293363293-363 K are compared with the empirical relation from Equation 1. The measured number densities follow the empirical vapor-pressure relation closely over the entire temperature range, with deviations well within the experimental uncertainties. Although absorption data were acquired over a broad temperature range at both probe wavelengths, only representative data points at selected temperatures are displayed in the figure to maintain visual clarity.
Figure 6: Rubidium vapor number density as a function of temperature for the MEMS vapor cell. Experimental number densities extracted from absorption measurements at 780 nm (red diamonds) and fluorescence detection at 420 nm (blue diamonds) are compared with the empirical vapor pressure model (dashed line). The vertical dashed line indicates the phase transition temperature at 312.45 K. Insets (a), (b), and (c) show magnified views around 298 K, 313 K, and 332 K, respectively.
The successful reproduction of the empirical vapor-pressure curve using both standard ($100$ mm) and MEMS ($2$ mm) vapor cells demonstrates the scalability of the absorption-based number density determination method. This result is significant for precision spectroscopy and the development of compact, chip-scale atomic devices, where accurate control of vapor density is crucial for optimal performance.
V Conclusion
We have presented a quantitative method for extracting the absolute atomic number density of warm rubidium vapor in chip-scale MEMS vapor cells using single-pass absorption spectroscopy. The approach is based on a multi-level Lindblad master equation formalism that incorporates hyperfine structure, Doppler averaging, optical pumping, and transit-time relaxation. All experimentally measured parameters, including vapor-cell temperature, optical path length, beam diameter, and optical power, are directly included in the model.
The method was experimentally validated using a MEMS vapor cell over a broad temperature range. Despite the intrinsically low optical depth associated with such short path lengths, the extracted number densities exhibit agreement with the empirical vapor-pressure relation of Alcock et al.. Consistent density values were obtained from absorption measurements on both the strong $5S_{1/2}\!\rightarrow\!5P_{3/2}$ (780.24 nm) transition and the weak $5S_{1/2}\!\rightarrow\!6P_{3/2}$ (420.29 nm) transition shows the applicability of the technique in weak-absorption regimes relevant to MEMS-scale devices.
The intrinsic photodetector dark current was used as a zero-transmission reference to provide a stable, noninvasive calibration method well suited to MEMS vapor cells, where complete absorption cannot be achieved without impractically high temperatures. The resulting fits reproduce the measured spectra with high fidelity ($R^{2}>0.99$). The demonstrated framework enables precise and reproducible characterization of vapor density in MEMS alkali cells.
The theoretical model described in this work may be extended to predict absorption spectra and to calculate the atomic number density of other atomic or molecular vapors, provided that the energy-level diagrams are known. Future work could also include additional transitions in the Rb atom to establish a complete formalism for the saturation absorption spectroscopy. These extensions will enhance the model’s versatility in atomic physics applications and the development of quantum technology.
Author Contributions
Sumit Achar and Shivam Sinha contributed equally to this work as co-first authors, theoretical model development, experimental design, data analysis, and manuscript writing. Ezhilarasan M. and Chandankumar R.: experimentation, data collection and analysis support. Arijit Sharma: conceptualization of the project, manuscript writing, review, editing, and overall project supervision.
Acknowledgments
Sumit Achar gratefully acknowledges financial support from the Council of Scientific & Industrial Research (CSIR, Govt. of India) through a Senior Research Fellowship (JRF). Shivam Sinha gratefully acknowledges the financial assistance provided by IIT Tirupati, facilitated through the half-time research assistantship (HTRA). Arijit Sharma acknowledges financial support from IIT Tirupati through the CAMOST grant and ISRO RESPOND project number RES-URSC-2022-005. Arijit Sharma thanks Dr. M. S. Giridhar, Dr. Jiju John and Dr. S. P. Karanth from the LEOS-ISRO (Laboratory for Electro-Optic Systems - Indian Space Research Organisation), Bengaluru, for sharing a sample MEMS rubidium (Rb) vapor cell for spectroscopic quantification through the ISRO RESPOND project.
Conflicts of Interest
The authors declare no conflicts of interest.
Appendix A Characterization of number density for 100 mm Rb vapor cell
We also determine the rubidium vapor number density in a 100 mm long vapor cell using absorption spectroscopy and quantitative spectral fitting. This analysis establishes a reliable, self-consistent characterization of the rubidium vapor number density in the 100 mm cell. The experimental schematic for the single-pass absorption spectroscopy is shown in Figure 7.
Figure 7: Schematic of the experimental setup for absolute absorption spectroscopy and frequency calibration for a 100 mm long vapor cell.
Figure 8 shows the measured absorption spectra as a function of laser detuning at three cell temperatures: $293.45~\pm~0.3$ K, $312.65~\pm~0.8$ K, and $339.65~\pm~3.3$ K. We fit the experimental spectra using a model that includes Doppler broadening, natural linewidth, and temperature-dependent atomic number density. The fitted spectra accurately reproduce the peak absorption, relative transition strengths, and off-resonant wings at all temperatures. The residuals (experiment minus fit), shown in the lower panels of Figure 8, remain within $\pm 0.25$ in absolute transmission.
Figure 8: Experimental and fitted absorption spectra as a function of detuning for three different temperatures: 293.45 K (red solid line for fit, blue dots for experiment), 312.65 K (black dashed line for fit, orange dots for experiment), and 339.65 K (blue dotted line for fit, green dots for experiment). The lower panels show the corresponding residuals (experimental minus fitted). The fitted spectra exhibit excellent agreement with experimental data across all temperatures, with residuals remaining within $\pm~0.25$ in absolute transmission.
From these fits, we extract the rubidium vapor number density as a function of temperature. Figure 10 shows the extracted number densities obtained independently from absorption measurements at 780 nm and 420 nm, together with the empirical rubidium vapor pressure model. The two experimental datasets show strong mutual agreement and closely follow the model over the full temperature range. The vertical dashed line at 312.45 K marks the solid–liquid phase transition of rubidium, where the temperature dependence of the number density changes slope, consistent with thermodynamic expectations. The insets highlight the agreement between experiment and model at selected temperatures.
Figure 9: Comparison of simulated and experimental transmission spectra of the 100 mm vapor cell at 310.65 K and 331.75 K. Solid lines represent the simulated spectra, and markers show the corresponding experimental data. The lower panels display the absolute residuals (experiment – simulation) for each temperature.
We further validate the extracted number densities by directly comparing simulated and experimental transmission spectra. Figure 9 presents this comparison at 310.65 K and 331.75 K using the corresponding fitted densities. The simulated spectra reproduce both the magnitude and spectral shape of the measured transmission. The residuals remain small across the full frequency scan, confirming that the extracted number densities provide a quantitatively accurate description of the rubidium vapor in the 100 mm cell.
Figure 10: Rubidium vapor number density as a function of temperature for a 100 mm vapor cell. Experimental number densities extracted from absorption measurements at 780 nm (red circles) and 420 nm (blue diamonds) are compared with the empirical vapor pressure model (dashed line). The vertical dashed line marks the phase transition point at 312.45 K. Insets (a), (b), and (c) show enlarged views of the data around 293 K, 313 K, and 341 K, respectively.
This analysis establishes a reliable, self-consistent characterization of the rubidium vapor number density in the 100 mm cell, providing a robust basis for subsequent precision spectroscopy and temperature-dependent studies.
To further validate the generality of the model, absorption spectroscopy was also performed on isotopically enriched rubidium vapor cells: an $86.5\%$ enriched 85Rb cell and a $93.6\%$ enriched 87Rb cell, each with a length of $100$ mm. Spectra were recorded at the same three cell temperatures ($293.45~\pm~0.3$ K, $312.65~\pm~0.8$ K, and $339.65~\pm~3.3$ K). For both isotopic compositions, the experimental spectra were well reproduced by the theoretical model, yielding $R^{2}>0.99$ in all cases. This consistency highlights the model’s versatility and reliability across different isotopic mixtures.
Although the NTC temperature sensors used in our experiment setup have an intrinsic accuracy of $\pm~0.3$ K, the observed error bars in the number density of the 100 mm long vapor cell at higher temperatures are larger than at lower temperatures. At higher temperatures, we observe that rubidium atoms in the vapor cell tend to deposit on the inner walls through which the laser beam passes. This introduces abrupt non-linearity in the absorption signal. To mitigate this issue, we introduced a slight temperature gradient along the cell, so that the flat glass window through which the laser beam passes remains slightly hotter than the center cylindrical region. Although this configuration helps us suppress the condensation on the flat window, it marginally increases the uncertainty in the temperature measurement.
Appendix B Absorption spectra at high power
[…]
At low probe powers, both the Beer-Lambert law and the iterative susceptibility approach yield virtually identical results, as expected in the linear regime, where intensity-dependent effects are negligible. However, at higher probe powers approaching or exceeding the saturation intensity, the Beer-Lambert law becomes inadequate, since the absorption depends on the local beam intensity, which decreases along the propagation axis.
We used an iterative approach that accounts for intensity-dependent susceptibility along the propagation path to accurately model absorption at higher intensities. Rather than treating the vapor cell as a uniform medium, we divided the 100 mm cell into $n$ equal slices along the laser propagation axis, where each slice can be considered to have constant incident laser intensity along its length. The number of slices $n$ can be adjusted based on computational resources and desired accuracy, with more slices providing higher precision at the cost of increased computation time.
The intensity variation through each slice can be expressed as [47],
$\frac{dI}{dz}=2k*\text{Im}\left\{\sqrt{1+\chi(I(z),\Delta,T)}\right\}I(z),$ (14)
where $\chi(I(z),\Delta,T)$ is the electric susceptibility, which mainly depends on the temperature (T) of the vapor cell, the detuning ($\Delta$) of the laser, and the local beam intensity ($I(z)$). This intensity dependence captures the saturation effects that become prominent at higher powers, allowing accurate modeling of the nonlinear absorption process beyond the weak-probe limit.
[…]
The strong agreement across different probe powers, from the linear weak-probe regime to the onset of saturation, confirms the reliability and versatility of our model. This validation supports its use for extracting accurate number density of alkali vapors and transition parameters beyond the weak-probe regime.
Appendix C Measurement of laser beam diameter
[…]
Figure 12: Beam profile analysis of the laser used in the experiment, measured using a CCD camera. The green and blue line shows the minor axis and major axis, respectively; the blue dot represents the centroid, and the white illustrates the $1/e^{2}$ elliptical beam diameter of the laser beam. This characterization is consistent with a distorted (elliptical) Gaussian beam and follows the ISO 11146 standard [48] for laser beam profiling.
The recorded beam profile was analyzed according to the ISO 11146 standard [48], which prescribes the use of second-order statistical moments to determine the beam width and orientation. Specifically, first-order moments were used to identify the centroid of the beam (center of mass), while second-order moments provided information about the spatial variance and the azimuthal tilt of the beam [49]. The beam diameters corresponding to the $1/e^{2}$ intensity level were extracted along the principal axes of the elliptical profile, which were rotated with respect to the CCD pixel grid. These diameters, initially calculated in pixel units, were converted to physical dimensions using the known pixel size. Based on this analysis, the beam diameter along the major axis was measured to be $2.22~\pm~0.04$ mm, while the minor axis diameter was found to be $1.52~\pm~0.03$ mm with the rotation angle of the principal axis 4.15.
Appendix D Measurement of transmitted power after commercial vapor cell
The transmitted laser intensity through the rubidium vapor cell was measured as a function of the incident intensity using a calibrated power meter. Figure 13 compares the experimental data with theoretical predictions obtained from the Beer–Lambert law. In the theoretical model, attenuation due to the glass cell walls was independently measured in the off-resonant region and incorporated into the transmission calculation. For a cell temperature of 293.15 K, the measured transmission follows the theoretical expectation across the full range of incident intensities investigated. The linear correlation observed between incident ($I_{0}$) and transmitted ($I(z)$) intensities confirms that absorption remains in the weak-probe regime, where the Beer–Lambert formalism accurately describes the system. The excellent agreement between experiment and theory indicates that both the vapor absorption and the cell wall losses have been accounted for in the model, thereby validating the use of absolute absorption spectroscopy for extracting quantitative number density information under these conditions.
Figure 13: Comparison of experimental and theoretical transmitted laser intensity through the rubidium vapor cell of length 100 mm at a temperature of 293.15 K. The blue stars represent the measured transmitted intensities obtained with a calibrated power meter, while the red circles with connecting lines indicate the theoretical values calculated from our model using the Beer–Lambert law, including off-resonant glass-cell wall losses.